OpenStudy (anonymous):

cot(csc inverse sqrt 10) Find the exact value, express in radians will fan and medal

OpenStudy (anonymous):

express in radians right !! how about "express is euros" or "express is centigrade"

OpenStudy (anonymous):



OpenStudy (anonymous):

find \(a\) via pythagoras then you can take any trig ration you like

OpenStudy (anonymous):

express in radians? did someone actually write that? it is not an angle you are looking for, it is just a number

OpenStudy (anonymous):

For the a I got sqrt 9

OpenStudy (anonymous):

you should find that the missing side is 3

OpenStudy (anonymous):

yeah \(\sqrt 9=3\) right?

OpenStudy (anonymous):

Oh yeah

OpenStudy (anonymous):

so cotangent is \(\frac{3}{1}=3\)

OpenStudy (anonymous):

but it sure as hell isn't three radians, it is the number three

OpenStudy (anonymous):

tell your math teacher that "express in radians" business is just plain nonsense

OpenStudy (xapproachesinfinity):

hehe was slow lol

OpenStudy (anonymous):

Thats okay, thanks :)

OpenStudy (anonymous):

hmm \[\csc(\theta)=\sqrt{10}\iff \sin(\theta)=\frac{1}{\sqrt{10}}\]

OpenStudy (anonymous):

it is very unlikely that the sine of anything is the square root of ten

OpenStudy (anonymous):

@xapproachesinfinity you got your triangle labelled goofy root ten should be the hypotenuse

OpenStudy (xapproachesinfinity):

lol not thinking hehe

OpenStudy (xapproachesinfinity):

at some point i was thinking of sin inverse not csc inverse lol